OpenStudy (anonymous):

if y varies directly as the square of x, and y=925 when x=25 find y whenx=20

OpenStudy (anonymous):

Based on the info given , I infer : y=x^2+c, where c is the constant 925=25^2+c 925=625+c or c=300 so, y=(20)^2+300 y=400+300 y=700 Please check if I my answer is right

OpenStudy (tkhunny):

It is no good. Direct and Inverse variation cannot have such an additive constant. \(y = k\cdot x^{2}\) Now, substitute the given values and solve for k.